OpenStudy (anonymous):

How do you solve 2x^2-11x+5=0 using the quadratic formula ? Please include all steps

OpenStudy (wolf1728):

First you have to define a, b and c a=2 b=-11 c=5

OpenStudy (wolf1728):

The formula is x = (-b +-sqroot(b^2 - 4ac)) / 2a x = (--11 +-sqroot([-11^2] - 4*2*5)) / 2*2 x = (11 +-sqroot(121 - 40)) / 4 x = (11 +- 9) / 4 x1= 20/4 = 5 x2 = 2/4 = 2

OpenStudy (wolf1728):

****************************************************************** CORRECTION x2 = 2/4 = .5 *********************************************************************